OpenStudy (anonymous):

solve the formula for H. f = 2gh h = ?

TheSmartOne (thesmartone):

Divide both sides by 2g...

TheSmartOne (thesmartone):

\[\frac{ f }{ 2g }=\frac{ 2gh }{ 2g }=?\]

OpenStudy (igreen):

\(f = 2gh\) Divide 2g to both sides: \(\dfrac{f}{2g} = gh\)

TheSmartOne (thesmartone):

^^^^

OpenStudy (igreen):

My bad its \(\dfrac{f}{2g} = h\)

OpenStudy (anonymous):

completely confused :(

OpenStudy (igreen):

To solve for h, you have to get rid of the 2g that is being multiplied on \(h\). To get rid of the 2g, you have to divide it to both sides, giving you f/2g = h